Advertisements
Advertisements
प्रश्न
An urn contains four tickets marked with numbers 112, 121, 122, 222 and one ticket is drawn at random. Let Ai (i = 1, 2, 3) be the event that ith digit of the number of the ticket drawn is 1. Discuss the independence of the events A1, A2, and A3.
Advertisements
उत्तर
One ticket can be drawn out of 4 tickets in 4C1 = 4 ways.
∴ n(S) = 4
According to given information,
Let A1 be the event that 1st digit of the number of ticket is 1
A2 be the event that 2nd digit of the number of ticket is 1.
A3 be the event that 3rd digit of the number of ticket is 1.
∴ A1 = {112, 121, 122}, A2 = {112}, A3 = {121}
∴ `"P"("A"_1) = ("n"("A"_1))/("n"("S")) = 3/4`,
`"P"("A"_2) = ("n"("A"_2))/("n"("S")) = 1/4`,
`"P"("A"_3) = ("n"("A"_3))/("n"("S")) = 1/4`
`{:("P"("A"_1) "P"("A"_2) = 3/16),("P"("A"_2) "P"("A"_3) = 1/16),("P"("A"_1) "P"("A"_3) = 3/16):}}` ...(i)
A1 ∩ A2 = {112}, A2 ∩ A3 = Φ, A1 ∩ A3 = {121}
`{:("P"("A"_1 ∩ "A"_2) = ("n"("A"_1 ∩ "A"_2))/("n"("S")) = 1/4),("P"("A"_2 ∩ "A"_3) = 0),("P"("A"_1 ∩ "A"_3) = 1/4):}}` ...(ii)
∴ From (i) and (ii),
`{:("P"("A"_1)*"P"("A"_2) ≠ "P"("A"_1 ∩ "A"_2)),("P"("A"_2)*"P"("A"_3) ≠ "P"("A"_2 ∩ "A"_3)),("P"("A"_1)*"P"("A"_3) ≠ "P"("A"_1 ∩ "A"_3)):}}` ...(iii)
∴ A1, A2, A3 are not pairwise independent
For mutual independent of events A1, A2, A3 We require to have
P(A1 ∩ A2 ∩ A3) = P(A1) P(A2) P(A3)
and P(A1) P(A2) = P(A1 ∩ A2),
P(A2) P(A3) = P(A2 ∩ A3),
P(A1) P(A3) = P(A1 ∩ A3)
∴ From (iii),
A1, A2, A3 are not mutually independent.
APPEARS IN
संबंधित प्रश्न
A card from a pack of 52 playing cards is lost. From the remaining cards of the pack three cards are drawn at random (without replacement) and are found to be all spades. Find the probability of the lost card being a spade.
A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, exactly two students solve the problem?
One-shot is fired from each of the three guns. Let A, B, and C denote the events that the target is hit by the first, second and third guns respectively. assuming that A, B, and C are independent events and that P(A) = 0.5, P(B) = 0.6, and P(C) = 0.8, then find the probability that at least one hit is registered.
A bag contains 3 red and 5 white balls. Two balls are drawn at random one after the other without replacement. Find the probability that both the balls are white.
Solution: Let,
A : First ball drawn is white
B : second ball drawn in white.
P(A) = `square/square`
After drawing the first ball, without replacing it into the bag a second ball is drawn from the remaining `square` balls.
∴ P(B/A) = `square/square`
∴ P(Both balls are white) = P(A ∩ B)
`= "P"(square) * "P"(square)`
`= square * square`
= `square`
Solve the following:
Find the probability that a year selected will have 53 Wednesdays
Two dice are thrown together. Let A be the event ‘getting 6 on the first die’ and B be the event ‘getting 2 on the second die’. Are the events A and B independent?
If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.
Refer to Question 1 above. If the die were fair, determine whether or not the events A and B are independent.
The probability that at least one of the two events A and B occurs is 0.6. If A and B occur simultaneously with probability 0.3, evaluate `"P"(bar"A") + "P"(bar"B")`
A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("B"/"A")`
Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: P1P2
If A and B are two events such that P(B) = `3/5`, P(A|B) = `1/2` and P(A ∪ B) = `4/5`, then P(A) equals ______.
Two events E and F are independent. If P(E) = 0.3, P(E ∪ F) = 0.5, then P(E|F) – P(F|E) equals ______.
If A and B are independent events, then A′ and B′ are also independent
If A and B′ are independent events, then P(A' ∪ B) = 1 – P (A) P(B')
Let A and B be independent events P(A) = 0.3 and P(B) = 0.4. Find P(A ∩ B)
Let Bi(i = 1, 2, 3) be three independent events in a sample space. The probability that only B1 occur is α, only B2 occurs is β and only B3 occurs is γ. Let p be the probability that none of the events Bi occurs and these 4 probabilities satisfy the equations (α – 2β)p = αβ and (β – 3γ) = 2βy (All the probabilities are assumed to lie in the interval (0, 1)). Then `("P"("B"_1))/("P"("B"_3))` is equal to ______.
Five fair coins are tossed simultaneously. The probability of the events that at least one head comes up is ______.
A die is thrown once. If E is the event that the number appearing is a multiple of 3, which set represents E?
A die is thrown once. If F is the event that the number appearing is even, which set represents F?
